OpenStudy (anonymous):

A jet can travel 4550 km in 5 hours with the wind. The return trip against the wind takes 6.5 hours. Find the rate of the jet with no wind.

OpenStudy (anonymous):

v=d/t With the wind, the jet is going 4550km/5hr=910km/hr Without the wind, the jet is going 4550km/6.5hr=700km/hr 700+w=910-w 2w=210km/hr w=105km/hr